OpenStudy (anonymous):

What is the area of a triangle with a base of 15 inches and a height of 4 inches?

OpenStudy (anonymous):

is it 30?

OpenStudy (arbershabani97):

a=1/2(b*h)

OpenStudy (aggie12341):

15*4=60

OpenStudy (anonymous):

what about 1/2 the base?

OpenStudy (aggie12341):

1/2 (15*4)=30

OpenStudy (anonymous):

so yeh its 30.

OpenStudy (arbershabani97):

yes

OpenStudy (aggie12341):

medal and ill give one back!